OpenStudy (specek18):

Using the inspection method, what constant term must the numerator of the rational expression x2+8x+11x+6 have so (x + 6) is one of its factors? For the numerator to have (x + 6) as a factor, its constant term must be

OpenStudy (specek18):

here is my next question...not multiple choice, will give medal

OpenStudy (aum):

If (x + 6) is a factor it implies if we put x = -6, f(x) should be zero. f(x) = x^2+8x+11x+6 f(-6) = ?

OpenStudy (specek18):

so now, we are looking for f....what's the best way to find it?

OpenStudy (aum):

Can you post a link of this question as well? I am wondering why they have it as 8x + 11x instead of 19x.

OpenStudy (specek18):

can you see that okay?

OpenStudy (specek18):

that would make a lot of sense....sorry.

OpenStudy (aum):

Yeah, I think you made a mistake in typing the problem. The rational expression is x^2 + 8x + 11 NOT x^2 + 8x + 11x + 6

OpenStudy (aum):

f(x) = x^2 + 8x + 11 If (x + 6) is a factor, then putting x = -6 should make f(x) = 0 f(-6) = (-6)^2 + 8(-6) + 11 = 36 - 48 + 11 = -1 We are left with -1 instead of zero. Therefore, we need to increase f(x) by 1 to get f(-6) to equal zero. So f(x) should be x^2 + 8x + 12 So the CONSTANT term should be 12 in order for (x + 6) to be a factor.

OpenStudy (specek18):

sorry i just wrote all of that out because i want to actually learn something:) haha thank you. i do understand how you got that answer now. would you be able to keep helping me?

OpenStudy (specek18):

i could put the new problem up sothat i can give you another medal:)

OpenStudy (specek18):

@aum

OpenStudy (aum):

close this one. post a new one and tag me and I will try to help.

OpenStudy (specek18):

:) thank you